# Complements of discriminants of real parabolic function singularities. II

### v2

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularity
Fig. 7. This diagram does not contain edges of multiplicity 2-2. Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2. Therefore, the virtual component associated with our polynomial is the unique remaining virtual component with $\mbox{Ind}=-3$. $\Box$
@@
It is easy to see that the polynomials (58) with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of components with $\mbox{Card}=336$ and $94$, respectively.
@@
Theorem 44.
There exists a $S(3)$-invariant polynomial $f\in\Xi_{2}$ realizing a virtual component with $\mbox{\rm Card}=60.$
Proof. One of the two virtual components, each of which has $\mbox{Card}=60$, contains a virtual function that models a generic polynomial $f_{0}$ of class $\Xi_{2}$ such that
1) all of its critical points are real,
2) the three highest critical values are positive and the other five values are negative,
3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is shown in (59),
4) the Morse indices of the points 1, 2, 3, and 5 are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
According to Proposition 32, there is a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 11 of [22], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), all odd Morse indices are equal to 1, and all even Morse indices are equal to 2.
Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by their increase, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga map (in the form of [10]) we can lift the path $i$ uniquely into a path $I:[0,1]\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 7 of [22], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [10], this is guaranteed by the fact that all quadratic forms corresponding to the limit critical values are elliptic. The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with value $c_{-}$ and three critical points with value $c_{+}$.
@@
Proof. The critical points of the polynomial $I(1)$ are ordered according to the previous ordering of the critical points of the polynomial $f_{0}$, from which they were obtained via continuous deformation along the path $I$. This ordering can be uniquely extended to a continuous system of orderings of the critical points of each polynomial $\tilde{f}$ from a sufficiently small neighborhood $U$ of the point $I(1)$ in $\Xi_{2}$. Define the map $\Lambda:U\to{\mathbb{R}}^{8}$, mapping each point $\tilde{f}\in U$ to the sequence of its correspondingly ordered critical values. By the versality property, this map is a local diffeomorphism. Let $\varepsilon>0$ be a very small number. For any permutation $\sigma\equiv\binom{1\ 2\ 3}{i_{1}\,i_{2}\,i_{3}}$ of critical points of $I(1)$ with critical value $c_{-}$, there is a unique point $\tilde{f}_{\sigma}\in U$, whose critical values at the critical points with numbers $i_{1},4,5,$ and $5+i_{1}$ are the same as those of $I(1)$, the values at the points $i_{2}$ and $5+i_{2}$ are equal to $c_{-}+\varepsilon$ and $c_{+}+\varepsilon$, and the values at the points $i_{3}$ and $5+i_{3}$ are equal to $c_{-}-\varepsilon$ and $c_{+}-\varepsilon$, respectively. Different permutations $\sigma$ of critical points define six different points $\tilde{f}_{\sigma}\in U$. All of these points are associated with the same virtual function, because any permutation $\sigma$ of the first three rows and columns in (59) and the simultaneous permutation of the last three rows and columns by the same rule $\sigma$ preserves the matrix.
@@
As in the proof of Proposition 33 (repeating the proof of Proposition 8), the six points $\tilde{f}_{\sigma}\in U$ are all mapped to each other by the action of elements of the group $S(3)$. Therefore, each such element maps a point from an arbitrarily small neighborhood $U$ of the polynomial $I(1)$ to another point from the same neighborhood. By continuity, it also maps the polynomial $I(1)$ to itself. $\Box$
@@
By the construction, the Card invariant of the virtual function associated with arbitrarily small perturbations of the polynomial $I(1)$ is equal to 60. This proves Theorem 44. $\Box$

### v3

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Fig. 7. This graph does not contain edges of multiplicities other than 1 and 1-1. Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2. Therefore, the virtual component associated with our polynomial is the only remaining virtual component with $\mbox{Ind}=-3$. $\Box$
@@
It is easy to check that the polynomials fCf-C with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of some components with $\mbox{Card}=336$ and $94$, respectively.
@@
Proposition 45.
There exists a $S(3)$-invariant polynomial $f\in\Xi_{2}$ that realizes a virtual component with $\mbox{\rm Card}=60.$
Proof. According to Proposition 41, two virtual components exist with $\mbox{Card}=60$. Our program shows that one of these components contains a virtual function that models a generic polynomial of class $\Xi_{2}$ such that
1) all of its critical points are real,
2) the three highest critical values are positive and the other five critical values are negative,
12536748
Figure 8. Coxeter-Dynkin graph for (59
3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is as follows:
@@
and hence its Coxeter-Dynkin graph is shown in Fig. 8,
4) the Morse indices of the points number 1, 2, 3, and 5 with respect to this order are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
According to Proposition 33, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 41 of [28], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2.
Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 7 of [28], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [12], this is guaranteed by the fact that all intersection forms corresponding to the limit positions $c_{-}$ and $c_{+}$ of the critical values of the polynomials $I(t)$ are elliptic. (These forms are given respectively by the upper left and the bottom right $3\times 3$ corners of the matrix (59).) The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with critical value $c_{-}$ and three critical points with critical value $c_{+}$.
@@
Proof. The critical points of the polynomial $I(1)$ are ordered according to the previous ordering of the critical points of the polynomial $f_{0}$, from which they were obtained via continuous deformation along the path $I$. This ordering can be uniquely extended to a continuous system of orderings of the critical points of each polynomial $\tilde{f}$ from a sufficiently small neighborhood $U$ of the point $I(1)$ in $\Xi_{2}$. Define the map $\Lambda:U\to{\mathbb{R}}^{8}$, mapping each point $\tilde{f}\in U$ to the sequence of its correspondingly ordered critical values. By the versality property, this map is a local diffeomorphism. Let $\varepsilon>0$ be a very small number. For any permutation $\sigma\equiv\binom{1\ 2\ 3}{i_{1}\,i_{2}\,i_{3}}$ of critical points of $I(1)$ with critical value $c_{-}$, there is a unique point $\tilde{f}_{\sigma}\in U$, whose critical values at the critical points with numbers $i_{1},4,5,$ and $5+i_{1}$ are the same as those of the polynomial $I(1)$, the values at the points $i_{2}$ and $5+i_{2}$ are equal respectively to $c_{-}+\varepsilon$ and $c_{+}+\varepsilon$, and the values at the points $i_{3}$ and $5+i_{3}$ are equal respectively to $c_{-}-\varepsilon$ and $c_{+}-\varepsilon$. Different permutations $\sigma$ define six different points $\tilde{f}_{\sigma}\in U$. These points are all associated with the same virtual function, because any permutation $\sigma$ of the first three rows and columns in (59) and the simultaneous permutation of the last three rows and columns by the same rule $\sigma$ preserves the matrix.
@@
As in the proof of Proposition 34, the six points $\tilde{f}_{\sigma}\in U$ are all mapped to each other by the action of elements of the group $S(3)$. The order of this group is also six, therefore each element of the group only permutes these six points. In particular, it maps a point from an arbitrarily small neighborhood $U$ of the polynomial $I(1)$ to another point in the same neighborhood. By continuity, it also maps the polynomial $I(1)$ to itself. $\Box$
@@
Due to the construction of the polynomial $I(1)$, the Card invariant of arbitrarily small generic perturbations of this polynomial equal to that of the initial polynomial $f_{0}$, that is, it is equal to 60. This proves Proposition 45. $\Box$